Two people, 27 km apart, setting out at the same time are together in 9 hours if they walk in the same direction but in 3 hours if they walk in opposite directions. Find each of their speeds.
Answer by jorel555(1290)

Let the 1st person be m, 2nd person be n. Walking in opposite direction, they cover 27 km in 3 hours, so m+n=27/3=9 km hr.
m+n=9
m=9-n
Walking in the same direction:
9m+27=9n
So:
9(9-n)+27=9n
81-9n+27=9n
18n=108
n=6
9-n=3
The 2nd person walks 6 km/h, the first walks 3 km/h.!!!!!!!!!!!!!!